# Constructing embedded surfaces for cellular embeddings of leveled spatial graphs

### v1

@@ 4.3 The algorithm for the surface construction
Consider a tree $T$ and the bipartite partition $X_{1},X_{2}$ of the vertices of $T$. To every vertex $v_{1}^{i}\in X_{1}$ assign a leveled spatial graph $\mathcal{G}_{i}$, $1\leq i\leq|X_{1}|$. The $\mathcal{G}_{i}$ are called blocks of the embedding. To every vertex $v_{2}^{j}\in X_{2}$, $1\leq j\leq|X_{2}|$, assign a connected planar graph $H_{j}$ with at least $\deg(v_{2}^{j})$ vertices. For each vertex $v_{1}^{i}$ adjacent to $v_{2}^{j}$, identify some vertices of $H_{j}$ with some vertices of the block $\mathcal{G}_{i}$ such that identified vertices belong to exactly one planar graph and one block. (See fig. 10)
Figure 10: A multi-leveled embedding with three blocks and two connected planar graphs. On the top left, the tree associated to the embedding.
@@ 4.3.1 Results of the algorithm
The algorithm gives a sufficient condition for a leveled spatial graph to be cellular embedded in an oriented surface. To investigate in how far the successful termination of the algorithm is also a necessary condition, we tried to construct a leveled spatial graph $\mathcal{G}$ for which the algorithm ends unsuccessfully for all possible choices of the spine. We have not been able to construct such an example. In all graphs we studied, it was always possible to find a spine for which theorem 1, theorem 2, proposition 1, or corollary 2 apply. As an example, consider the spatial graph $\mathcal{G}$ described in example 1. Although for the given spine the algorithm does not find a surface $\mathcal{S}$ where $\mathcal{G}$ cellular embeds, it is possible to choose a non-hamiltonian spine for $\mathcal{G}$, shown in fig. 12, for which algorithm 2 terminates successfully. Furthermore, a different hamiltonian spine for $\mathcal{G}$ can be chosen, shown in fig. 13, which yields a leveled embedding with five levels. Corollary 2 guarantees the existence of a surface where $\mathcal{G}$ cellular embeds, and indeed algorithm 1 terminates successfully if applied to this leveled embedding. Note that it follows from this example that all hamiltonian leveled spatial graphs with a spine list of the form $(1^{1},1^{2},1^{3},2^{1},2^{2},2^{3},\dots,n^{1},n^{2},n^{3},1^{3},1^{2},1^{1},2^{3},2^{2},2^{1},\dots,n^{3},n^{2},n^{1})$ for $n\in\mathbb{N}$ can be cellular embedded by applying algorithm 1 to a different choice of the spine, analogously to the example.
@@
It is important to highlight that admitting a leveled embedding for a spatial graph $\mathcal{G}$ is not a necessary condition for the existence of a surface $\mathcal{S}$ where $\mathcal{G}$ cellular embeds. Indeed, the trefoil knot with an unknotting tunnel is an example of a spatial graph that does not admit a leveled embedding, because for every unknotted cycle chosen as spine, the fragment will cross the spine non-trivially. It does though have a cellular embedding in a torus, as shown in fig. 14.
@@
In [1], it is shown that any spatial graph $\mathcal{G}$ can be embedded in a 3-book, where the vertices lie on the back of the 3-book and edges may be contained in several leaves (i.e. the pages of the 3-book). Placing a vertex on every edge of $\mathcal{G}$ which intersects the back of the 3-book and adding an edge between two consecutive vertices on the back which are not already adjacent yields a leveled embedding of a spatial graph $\mathcal{G}^{\prime}$ in three levels such that $\mathcal{G}\subset\mathcal{G}^{\prime}$. In other words, every spatial graph has a super-graph with a leveled embedding for which algorithm 2 ends successfully.
@@ 5 Outlook
A similar conjecture can be stated for hamiltonian leveled spatial graphs. The situation shown in fig. 13 is representative for all examples of hamiltonian leveled spatial graphs that we studied.
@@
Cellular embeddings of graphs are related to circular embeddings. A circular embedding is a cellular embedding $(\mathcal{G},\mathcal{S})$ where the boundaries of the faces of $\mathcal{S}\setminus\mathcal{G}$ are cycles of $\mathcal{G}$. A famous open conjecture in graph theory states that every 2-connected graph has a circular embedding [6]. In case of a 2-connected hamiltonian leveled spatial graph $\mathcal{G}$ for which algorithm 1 ends successfully, it is possible to determine whether the resulting cellular embedding is circular by applying 1 to all faces of the cellular embedding: if every edge of $\mathcal{G}$ appears in the boundary of exactly two faces, then the embedding is circular. The embedding $(\mathcal{G},\mathcal{S})$ is ensured to be circular if no cylinder attachments merge disks in the construction of $\mathcal{S}$ as performed in algorithm 1.

### v2

@@ 4.3 The algorithm for the surface construction
Consider a tree $T$ and the bipartite partition $X_{1},X_{2}$ of the vertices of $T$. To every vertex $v_{1}^{i}\in X_{1}$ assign a leveled spatial graph $\mathcal{G}_{i}$, $1\leq i\leq|X_{1}|$. The $\mathcal{G}_{i}$ are called blocks of the embedding. To every vertex $v_{2}^{j}\in X_{2}$, $1\leq j\leq|X_{2}|$, assign a connected planar graph $H_{j}$ with at least $\deg(v_{2}^{j})$ vertices. For each vertex $v_{1}^{i}$ adjacent to $v_{2}^{j}$, identify some vertices of $H_{j}$ with some vertices of the block $\mathcal{G}_{i}$ such that identified vertices belong to exactly one planar graph and one block. (See Figure 11 and Figure 12)
Figure 11: A multi-leveled embedding with three blocks and two connected planar graphs.
Figure 12: The tree associated to the embedding of Figure 11.
@@ 4.3.1 Results of the algorithm
The algorithm gives a sufficient condition for a leveled spatial graph to be cellular embedded in an oriented surface. To investigate in how far the successful termination of the algorithm is also a necessary condition, we tried to construct a leveled spatial graph $\mathcal{G}$ for which the algorithm ends unsuccessfully for all possible choices of the spine. We have not been able to construct such an example. In all graphs we studied, it was always possible to find a spine for which Theorem 1, Theorem 2, Proposition 1, or Corollary 2 apply. As an example, consider the spatial graph $\mathcal{G}$ described in Example 1. Although for the given spine the algorithm does not find a surface $\mathcal{S}$ where $\mathcal{G}$ cellular embeds, it is possible to choose a non-hamiltonian spine for $\mathcal{G}$, shown in Figure 14, for which Algorithm 2 terminates successfully. Furthermore, a different hamiltonian spine for $\mathcal{G}$ can be chosen, shown in Figure 15, which yields a leveled embedding with five levels. Corollary 2 guarantees the existence of a surface where $\mathcal{G}$ cellular embeds, and indeed Algorithm 1 terminates successfully if applied to this leveled embedding. Note that it follows from this example that all hamiltonian leveled spatial graphs with a spine list of the form $(1^{1},1^{2},1^{3},2^{1},2^{2},2^{3},\dots,n^{1},n^{2},n^{3},1^{3},1^{2},1^{1},2^{3},2^{2},2^{1},\dots,n^{3},n^{2},n^{1})$, for $n\in\mathbb{N}$, can be cellular embedded by applying Algorithm 1 to a different choice of the spine, analogously to the example.
@@
It is important to highlight that admitting a leveled embedding for a spatial graph $\mathcal{G}$ is not a necessary condition for the existence of a surface $\mathcal{S}$ where $\mathcal{G}$ cellular embeds. Indeed, the trefoil knot with an unknotting tunnel is an example of a spatial graph that does not admit a leveled embedding, because for every unknotted cycle chosen as spine, the fragment will cross the spine non-trivially. It does though have a cellular embedding in a torus, as shown in Figure 16.
@@
In Atn (68), it is shown that any spatial graph $\mathcal{G}$ can be embedded in a 3-book, where the vertices lie on the back of the 3-book and edges may be contained in several leaves (i.e. the pages of the 3-book). Placing a vertex on every edge of $\mathcal{G}$ which intersects the back of the 3-book and adding an edge between two consecutive vertices on the back which are not already adjacent yields a leveled embedding of a spatial graph $\mathcal{G}^{\prime}$ in three levels such that $\mathcal{G}\subset\mathcal{G}^{\prime}$. In other words, every spatial graph has a super-graph with a leveled embedding for which Algorithm 2 ends successfully.
@@ 5 Outlook
A similar conjecture can be stated for hamiltonian leveled spatial graphs. The situation shown in Figure 15 is representative for all examples of hamiltonian leveled spatial graphs that we studied.
@@
Cellular embeddings of graphs are related to circular embeddings. A circular embedding is a cellular embedding $(\mathcal{G},\mathcal{S})$ where the boundaries of the faces of $\mathcal{S}\setminus\mathcal{G}$ are cycles of $\mathcal{G}$. A famous open conjecture in graph theory states that every 2-connected graph has a circular embedding Jae (85). In case of a 2-connected hamiltonian leveled spatial graph $\mathcal{G}$ for which Algorithm 1 ends successfully, it is possible to determine whether the resulting cellular embedding is circular by applying 1 to all faces of the cellular embedding: if every edge of $\mathcal{G}$ appears in the boundary of exactly two faces, then the embedding is circular. The embedding $(\mathcal{G},\mathcal{S})$ is ensured to be circular if no cylinder attachments merge disks in the construction of $\mathcal{S}$ as performed in Algorithm 1.